Styles ont Paragraph experiment to determine the most effective way to teach safety principles Question 8.8. An applied four different teaching methods. Some employees were given programmed instruction booklets and worked through the course at their own pace. Other employees attern A third group watched a television presentation, and a fourth group was divided into small discussion groups. A high of 10 was possible A sample of five tests was selected from each group. The test grade results were: At the 0.01 level, what is the critical value? (Points 2) ? 1.00 3.24 1.96 ? 529
media%2F2d8%2F2d8d214f-bee8-4cec-99a6-64
Answer #1

here degree of freedom for numerator =groups-1=4-1=3

degree of freedom for denminator =N-k=4*5-4=16

hence for (3,16 )df and 0.01 level ; critical value of F =5.29

18)

degree of freedom for treatments =number of treatments -1=3-1=2